# Evaluating alignment between humans and neural network representations in image-based learning tasks

### v2

@@ 1 Introduction
  • Of the tested human-aligned neural networks, only one of the three methods, gLocal [10], improved task accuracy compared to not-aligned baselines. This improvement in accuracy was accompanied by improved human alignment for some of the tested models.
@@ 4 Model-based analyses
We trained linear models to predict either reward or category membership from each neural network model’s extracted representations. The models were provided with image-target pairs until trial $t-1$ as training data and made predictions for the image on trial $t$. For the category learning task, we used an $\ell_{2}$ regularised logistic regression model, and for the reward learning task, we used a Bayesian linear regression model with spherical Gaussian priors (see Appendix A for other tested models and Appendix D for the corresponding results). We used the estimates from the linear models to predict participant choice using mixed-effects logistic regression in leave-one-trial-out cross-validation. For the category learning task, we regressed the probability estimates of the logistic regression models onto participant choice. For the reward learning task, we regressed the reward estimate differences between the left and the right options onto choice. Example learning curves for the two tasks are shown in Fig. 2C and Fig. 2D. Finally, we measure alignment to human choices using McFadden’s $R^{2}$ [57], which is computed as follows:
@@
We observed all but $3$ of the representations we tested can do our task and predict human behaviour above chance level across the two tasks (as visualized in Fig. 3A and Fig. 3B). CLIP models were the top $7$ ($6$) models for the category (reward) learning task in predicting human choices. In total, $16$ ($7$) of the $86$ candidate representations predicted participant behaviour better than the ground truth representations that were used to generate the task. Of these $16$ ($7$) representations, $14$ ($6$) were CLIP models. One was a large vision transformer, trained in a supervised manner on ImageNet [58]. A human-aligned variant of DINO-v2 provided a better fit than the generative task representations in the category learning task. The rest of the supervised and self-supervised vision models, as well as the language models, had a heterogeneous distribution in how well they predicted human behaviour. To provide better intuition for how human participants and CLIP were similar, we display example trials where both CLIP and humans make the same incorrect decisions in Fig. 13 in Appendix D.
@@
Figure 5: Lower intrinsic dimensionality is linked with higher alignment only for the multimodal models.
@@
First, we found that both Harmonization and DreamSim led the models to predict human choices worse compared to their non-aligned baseline alternatives (Fig. 7A). gLocal transformation improved how well representations predicted human choice for two models (large DINOv2 and ResNet) but not for the five other models tested (CLIP variants and base DINOv2). Alignment hurt task accuracy for all Harmonization and DreamSim models, whereas for gLocal, all but two models showed improved performance (Fig. 7B). Class separation was lower for all Harmonization and DreamSim models, whereas it increased for all gLocal models tested (Fig. 7C). We also observed that the similarity between the representations and the task embedding decreased after alignment for Harmonization and DreamSim, but it increased in most of the models after gLocal alignment (Fig. 7D). Lastly, we observed heterogeneous patterns in the change of intrinsic dimensionality across the three alignment methods, with gLocal reducing the intrinsic dimensionality for all but one of the tested methods (Fig. 7E).
Figure 7: We compared models aligned to humans through three different methods against baselines that had the same architecture and that were pretrained on the same data. Only two of the gLocal models showed increased alignment in our tasks. Here plotted are (A) McFadden’s $R^{2}$, (B) task accuracy, (C) class-separation, (D) similarity with the task embedding, and (D) intrinsic dimensionality across model sizes and loss functions.
@@
Lastly, to better characterise how our cognitive tasks fit in the alignment literature, we compared them to previously established measures (Fig. 8). We found the strongest correlation with the THINGS odd-one-out judgements [3, 27] ($\rho=0.54$ for zero-shot, and $\rho=0.61$ for probing). Given the two tasks use the same images, and the ground truth of our tasks was constructed from the odd-one-out judgements, this strong relationship is expected. However, the correlations are still moderate, indicating important differences across tasks. Comparisons with an independent similarity judgement [23] and a two-alternative forced choice [12] showed weaker correlations ($\rho=0.45$ and $\rho=0.28$, respectively). Lastly, we compared alignment in our task to alignment on the ClickMe dataset [76], which was used to build the Harmonization models [11]. We observed a negative correlation ($\rho=-0.43$) here, suggesting that pixel-level alignment and semantically bound global image alignment might be at odds.
@@ 5 Discussion
Figure 8: How do our tasks compare to other alignment methods? Our tasks offer similar (but not identical) results with similarity judgement tasks. There is a strong negative relationship with the ClickMe dataset, which focuses on localised pixel-level alignment.
@@ 5.1 Related work
Lastly, we found that two methods designed for increasing human alignment, Harmonization [11] and DreamSim [12], actually hurt alignment in our tasks. On the other hand, gLocal [10] improved both performance and human alignment for some models but not all of them. However, the gLocal transform heavily utilises the THINGS dataset, as it made use of the triplet odd-one-out similarity judgement data [27], making it difficult to interpret how well it generalises to other settings. Taken together, these results highlight the importance of studying how well different alignment methods transfer across tasks, as we have done in this work.
@@ Appendix A Methods
In addition to the models described above, we tested sparse models. For the category learning task, this was simply done by changing the penalty terms to consider the $\ell_{1}$ norm. For the reward learning task, the prior over-the-model weights were changed. Instead of using a spherical Gaussian, each weight’s prior had a different scale. Lastly, to account for the different number of features each representation had, we tested models trained on representations obtained through principal component analysis (PCA). For each representation and participant, we took the first $49$ principal components over all images shown to the participant. We selected a $49$ dimensional space because it matches the size of the embedding used to generate the task. None of the models described in this section predict human choice better than models reported in the main text. These results are shown in Table LABEL:tab:_all_results_table.
@@
Appendix B Robustness to the exclusion criteria and the choice of metric
Below in Fig. 9, we show the modelling results before any exclusion criteria are applied. These $R^{2}$ values are highly correlated with those after exclusion ($\rho=0.96$ for category learning, $\rho=0.66$ for reward learning), and our qualitative findings stay the same, where the top ranking models are CLIP variants. Similarly, if we consider % accuracy of models in predicting human choice as opposed to McFadden’s $R^{2}$, we get very similar results ($\rho=0.97$ for category learning, $\rho=0.93$ for reward learning), as shown in Fig. 10.
Figure 9: Model fits before applying any exclusion criteria for (A) category learning and (B reward learning
Figure 10: Accuracies of models in predicting human choices for (A) category learning and (B) reward learning
Appendix C Testing aligned models on other datasets
Below, we also report how different alignment methods perform on different datasets (Fig. 11). Harmonization is less human-like on the two external tested datasets compared to baselines. DreamSim shows mixed results for the Peterson et al. [23] dataset, but it shows improvement on the NIGHTS dataset [12]. This is not surprising, as this dataset was used to build DreamSim. Lastly, gLocal shows mixed results.
@@ Appendix C Testing aligned models on other datasets
Appendix D Additional results
Below we provide some additional results supporting our claims in the main text. Fig. 12 shows participants can do both tasks above chance level very early on in the task. Fig. 13 shows some incorrect choices made by humans and also by CLIP models, and Fig. 14 shows pairwise correlations between the factors we investigated that contribute to alignment. Additionally, Table LABEL:tab:_all_results_table shows negative log likelihoods for each tested representation across both tasks and across the tested linear models. Table LABEL:tab:alignment shows the factors impacting alignment for each set of representations.
@@ Appendix D Additional results
Table 1: Details on Models and Model Performance. Here, we report the model size, the type, and the negative log-likelihoods (NLL) for different tasks and different learning models. For a given representation and task, the best-performing learning model’s NLL is highlighted in bold. The models are ordered such that the ones reported higher have lower negative log likelihoods across the two asks, as measured through the performance of the linear models reported in the main text.
Category Learning NLL Reward Learning NLL
Model Name Type Parameters (Million) Main Sparse PCA Main Sparse PCA
CLIP-ViT-L/14 Multimodal 427.62 6556.57 7048.70 7350.88 2893.09 3293.56 3235.11
CLIP-ViT-B/16 Multimodal 149.62 6629.96 6996.63 7341.26 2896.44 3215.89 3229.61
CLIP-ViT-B/32 Multimodal 151.28 6660.77 7047.67 7363.74 2946.11 3197.59 3221.68
CLIP-RN50x4 Multimodal 178.3 6746.92 6930.88 7332.79 2932.43 3074.20 3053.85
OpenCLIP-ViT-H-Laion-2B Multimodal 986.11 6626.19 7063.46 7309.82 3056.02 3258.59 3179.22
CLIP-RN101 Multimodal 119.69 6802.21 7000.14 7307.13 2909.52 3047.88 3048.06
gLocal-CLIP-ViT-L/14 Multimodal 427.62 6673.51 6913.53 7376.33 3076.15 3132.30 3227.68
OpenCLIP-ViT-G-Laion-2B Multimodal 1366.68 6660.27 7148.54 7381.17 3089.96 3247.13 3221.05
ViT-H/14 Supervised 633.47 6746.31 7200.56 7256.95 3019.41 3211.69 3177.17
OpenCLIP-ViT-L-Laion-400M Multimodal 427.62 6711.45 7187.83 7309.94 3104.35 3297.51 3196.23
CLIP-RN50 Multimodal 102.01 6868.09 7065.86 7367.29 2950.57 3081.82 3080.26
OpenCLIP-ViT-B-Laion-400M Multimodal 151.28 6785.17 7143.46 7307.26 3103.29 3270.55 3204.41
gLocal-CLIP-RN50 Multimodal 102.01 6796.68 7002.53 7426.53 3123.73 3249.53 3180.56
SLIP-Large Multimodal 389.3 6841.71 7189.82 7409.62 3095.30 3325.74 3221.94
gLocal-DINO-V2-ViT-L Self-Supervised 304.37 6802.88 7084.55 7379.66 3156.79 3365.05 3228.71
Task Task 0.17 6976.13 6987.36 6876.38 3025.50 2970.23 2968.43
SLIP-Base Multimodal 170.62 6943.38 7225.39 7414.12 3083.30 3286.60 3231.93
CLIP-Base (SLIP) Multimodal 149.62 6994.92 7182.08 7431.45 3084.21 3326.63 3228.91
Universal Sentence Encoder Language 8.54 7080.84 7161.20 7404.60 3040.20 3126.09 3143.28
RoBERTa Language 124.65 7098.94 7197.73 7344.64 3035.65 3171.26 3178.74
SLIP-Small Multimodal 104.71 7023.78 7201.66 7303.53 3135.09 3282.03 3288.91
gLocal-OpenCLIP-ViT-B-Laion-400M Multimodal 151.28 6987.08 7187.30 7418.70 3178.26 3286.08 3205.86
CLIP-Large (SLIP) Multimodal 367.25 7009.00 7262.82 7433.40 3164.10 3450.09 3249.10
DINO Self-Supervised 23.51 7106.35 7262.30 7447.87 3094.62 3233.45 3165.34
SimCLR-Base (SLIP) Self-Supervised 106.8 7087.67 7257.07 7409.17 3114.10 3288.52 3325.40
fastText Language NA 7146.80 7299.53 7323.10 3066.87 3236.48 3180.89
gLocal-ResNet-50 Supervised 25.56 7075.16 7156.32 7439.82 3146.53 3286.20 3289.74
ConvNeXt-L Supervised 197.77 7120.20 7268.95 7387.76 3114.50 3203.61 3205.56
BERT Language 109.48 7184.10 7261.03 7422.32 3080.85 3169.49 3178.61
DreamSim-CLIP-ViT-B Multimodal 151.28 7110.21 7267.10 7423.03 3162.03 3277.41 3240.57
ResNet-18 Supervised 11.69 7183.59 7304.01 7370.80 3091.07 3194.43 3156.70
ResNet-34 Supervised 21.8 7205.51 7355.84 7378.90 3073.46 3225.14 3161.01
SimCLR-Large (SLIP) Self-Supervised 325.35 7042.58 7212.19 7475.49 3241.43 3406.53 3295.74
gLocal-ResNet-18 Supervised 11.69 7141.57 7236.79 7437.38 3155.14 3287.74 3386.47
ResNet-50 Supervised 25.56 7058.73 7249.82 7410.53 3244.52 3312.13 3180.09
CLIP-Small (SLIP) Multimodal 85.29 7073.67 7221.35 7344.83 3231.71 3370.81 3284.94
MoCoV2 Self-Supervised 23.51 7198.66 7262.55 7470.35 3112.54 3198.73 3160.91
ConvNeXt-B Supervised 88.59 7185.77 7282.16 7434.23 3128.76 3174.86 3167.18
Swin-T Supervised 28.29 7180.41 7417.50 7432.17 3156.21 3271.34 3254.49
Cornet-R Supervised 5.21 7179.79 7335.57 7370.35 3165.64 3219.80 3210.31
ConvNeXt-S Supervised 50.22 7214.27 7332.72 7464.86 3134.08 3215.19 3197.22
ada-002 Language 350∗∗ 7106.84 7264.09 7433.11 3243.41 3407.04 3253.14
Swin-S Supervised 49.61 7228.70 7423.84 7374.90 3136.99 3261.96 3270.30
VicReg Self-Supervised 23.51 7154.49 7270.59 7453.93 3219.45 3229.88 3188.65
Swin-B Supervised 87.77 7205.77 7343.31 7374.76 3168.50 3189.88 3261.50
Barlow Twins Self-Supervised 23.51 7156.34 7267.88 7455.53 3219.42 3239.21 3196.55
ViT-L/32 Supervised 306.54 7259.64 7404.67 7432.01 3151.54 3211.50 3271.45
DreamSim-OpenCLIP-ViT-B Multimodal 151.28 7190.91 7317.56 7437.87 3226.48 3332.33 3274.05
SwAV Self-Supervised 23.51 7133.95 7221.40 7479.26 3303.51 3277.18 3177.67
PiRL Self-Supervised 23.51 7133.95 7216.29 7482.89 3303.51 3277.18 3179.97
DistilBERT Language 66.36 7229.13 7288.39 7476.50 3220.91 3325.47 3308.40
ResNet-152 Supervised 60.19 7178.87 7316.28 7442.82 3276.84 3282.15 3222.57
ViT-B/16 Supervised 86.57 7258.40 7414.03 7455.65 3201.05 3274.24 3296.59
SimCLR Self-Supervised 23.51 7135.90 7221.14 7483.27 $3$.31 3268.50 3191.66
SimCLR-Small (SLIP) Self-Supervised 41.09 7138.05 7297.78 7380.27 3337.43 3377.40 3425.74
ResNet-101 Supervised 44.55 7136.56 7245.44 7405.02 3339.84 3292.09 3203.80
ConvNeXt-T Supervised 28.59 7324.95 7362.84 7454.36 3151.89 3239.63 3261.15
DINO-V2-ViT-L Self-Supervised 304.37 7209.51 7429.50 7412.64 3272.52 3473.91 3313.55
Jigsaw Self-Supervised 23.51 7360.46 7351.58 7511.62 3133.06 3209.14 3244.48
ViT-B/32 Supervised 88.22 7271.29 7466.39 7402.88 3225.13 3280.83 3285.52
DINO-V2-ViT-G Self-Supervised 1136.48 7211.42 7461.45 7437.81 3303.16 3421.66 3351.46
RotNet Self-Supervised 23.51 7343.46 7430.61 7496.19 3173.14 3210.07 3249.09
DINO-V2-ViT-B Self-Supervised 86.58 7259.44 7474.31 7391.35 3291.73 3365.58 3326.81
ViT-L/16 Supervised 304.33 7405.88 7470.41 7439.83 3156.49 3213.36 3283.52
CORnet-Z Supervised 2.07 7414.85 7432.33 7458.41 3160.30 3243.43 3231.07
DINO-V2-ViT-S Self-Supervised 22.06 7292.59 7448.97 7400.63 3354.21 3475.93 3301.97
Harmonization-ConvNeXt-T Supervised 28.59 7440.67 7460.92 7482.18 3221.63 3291.88 3284.43
gLocal-DINO-V2-ViT-B Self-Supervised 86.58 7400.35 7513.70 7538.13 3270.68 3393.64 3375.12
Harmonization-ViT-B/16 Supervised 86.57 7501.41 7498.24 7511.05 3291.06 3328.95 3342.78
Harmonization-MaxViT-T Supervised 30.96 7515.66 7524.84 7515.16 3309.47 3339.27 $3$.26
Cornet-S Supervised 53.42 7520.58 7522.31 7512.14 3307.13 3332.08 3336.97
Harmonization-RN50 Supervised 25.61 7508.76 7501.03 7496.49 3323.09 3314.34 3323.73
Harmonization-VGG16 Supervised 138.36 7501.40 7512.61 7522.21 3342.20 3440.21 3292.75
Harmonization-LeViT-S Supervised 8.88 7521.85 7520.23 7483.63 3327.08 3344.02 3337.67
Harmonization-EfficientNetB0 Supervised 5.33 7524.88 7523.05 7498.16 3338.08 3337.96 3321.88
Table 2: Investigated factors that contribute to alignment for each model
Model Name Task Accuracy Training Images (Millions) Class-Separation ($R^{2}$) CKA Intrinsic Dimensionality
CLIP-ViT-L/14 0.72 400 0.58 0.58 14.09
CLIP-ViT-B/16 0.72 400 0.60 0.61 14.90
CLIP-ViT-B/32 0.72 400 0.59 0.61 16.02
CLIP-RN50x4 0.71 400 0.58 0.60 15.91
OpenCLIP-ViT-H-Laion-2B 0.73 2000 0.54 0.53 14.90
CLIP-RN101 0.70 400 0.58 0.61 16.33
gLocal-CLIP-ViT-L/14 0.76 400 0.83 0.82 14.16
OpenCLIP-ViT-G-Laion-2B 0.73 2000 0.54 15.01
ViT-H/14 0.70 1.28 0.65 0.47 14.41
OpenCLIP-ViT-L-Laion-400M 0.72 400 0.54 0.53 15.56
CLIP-RN50 0.69 400 0.58 0.64 17.86
OpenCLIP-ViT-B-Laion-400M 0.71 400 0.55 0.57 17.14
gLocal-CLIP-RN50 0.74 400 0.84 0.79 17.06
SLIP-Large 0.70 15 0.44 0.48 20.82
gLocal-DINO-V2-ViT-L 0.74 142 0.85 0.75 11.76
Task 0.78 NA NA 1.00 NA
SLIP-Base 0.70 15 0.44 0.51 21.22
CLIP-Base (SLIP) 0.69 15 0.40 0.44 23.54
Universal Sentence Encoder 0.67 NA NA 0.44 NA
RoBERTa 0.68 NA NA 0.47 NA
SLIP-Small 0.69 15 0.45 0.51 21.24
gLocal-OpenCLIP-ViT-B-Laion-400M 0.71 400 0.77 0.56 17.12
CLIP-Large (SLIP) 0.69 15 0.39 0.41 23.47
DINO 0.66 1.28 0.41 0.35 21.48
SimCLR-Base (SLIP) 0.67 15.00 0.48 0.46 18.79
fastText 0.67 NA NA 0.35 NA
gLocal-ResNet-50 0.71 1.28 0.86 0.68 15.67
ConvNeXt-L 0.64 1.28 0.53 0.19 18.09
BERT 0.66 0.00 0.00 0.12 0.00
DreamSim-CLIP-ViT-B 0.68 400 0.53 0.48 16.64
ResNet-18 0.65 1.28 0.48 0.44 21.53
ResNet-34 0.65 1.28 0.49 0.44 19.91
SimCLR-Large (SLIP) 0.68 15.00 0.49 0.47 17.34
gLocal-ResNet-18 0.70 1.28 0.81 0.66 19.82
ResNet-50 0.66 1.28 0.50 0.43 18.72
CLIP-Small (SLIP) 0.67 15.00 0.41 0.44 22.65
MoCoV2 0.65 1.28 0.47 0.48 20.39
ConvNeXt-B 0.64 1.28 0.52 0.15 17.84
Swin-T 0.63 1.28 0.50 0.31 19.28
Cornet-R 0.64 1.28 0.50 0.49 20.70
ConvNeXt-S 0.63 1.28 0.49 0.11 17.94
ada-002 0.71 NA NA 0.57 NA
Swin-S 0.63 1.28 0.54 0.30 18.12
VicReg 0.65 1.28 0.39 0.36 20.35
Swin-B 0.63 1.28 0.55 0.29 18.13
Barlow Twins 0.65 1.28 0.39 0.36 20.35
ViT-L/32 0.61 1.28 0.44 0.29 22.48
DreamSim-OpenCLIP-ViT-B 0.67 400 0.51 0.44 16.90
SwAV 0.64 1.28 0.45 0.47 19.21
PiRL 0.64 1.28 0.45 0.47 19.21
DistilBERT 0.66 NA NA 0.21 NA
ResNet-152 0.65 1.28 0.52 0.37 18.43
ViT-B/16 0.62 1.28 0.47 0.27 19.78
SimCLR 0.64 1.28 0.45 0.47 19.21
SimCLR-Small (SLIP) 0.66 15.00 0.47 0.47 19.43
ResNet-101 0.65 1.28 0.52 0.39 17.97
ConvNeXt-T 0.63 1.28 0.50 0.19 19.00
DINO-V2-ViT-L 0.65 142.00 0.62 0.25 11.98
Jigsaw 0.60 1.28 0.37 0.22 23.31
ViT-B/32 0.61 1.28 0.44 0.27 20.90
DINO-V2-ViT-G 0.65 142.00 0.68 0.25 11.28
RotNet 0.62 1.28 0.38 0.14 27.08
DINO-V2-ViT-B 0.65 142.00 0.60 0.26 12.94
ViT-L/16 0.61 1.28 0.44 0.23 26.11
CORnet-Z 0.61 1.28 0.39 0.36 28.18
DINO-V2-ViT-S 0.64 142.00 0.57 0.27 14.24
Harmonization-ConvNeXt-T 0.58 1.28 0.32 0.08 22.64
gLocal-DINO-V2-ViT-B 0.63 142.00 0.80 0.22 12.85
Harmonization-ViT-B/16 0.54 1.28 0.36 0.12 16.42
Harmonization-MaxViT-T 0.52 1.28 0.16 0.05 4.86
Cornet-S 0.53 1.28 0.11 0.06 123.68
Harmonization-RN50 0.52 1.28 0.43 0.04 18.38
Harmonization-VGG16 0.54 1.28 0.31 0.07 31.88
Harmonization-LeViT-S 0.51 1.28 0.11 0.02 29.96
Harmonization-EfficientNetB0 0.51 1.28 0.08 0.00 10.54

### v3

@@ 1 Introduction
  • Of the tested human-aligned neural networks, no method consistently improved human alignment in our tasks compared to non-aligned baselines. However, two of the methods (Harmonization [11] and gLocal [10]) yielded improvements in task accuracy on average.
@@ 4 Model-based analyses
We trained linear models to predict either reward or category membership from each neural network model’s extracted representations. The models were provided with image-target pairs until trial $t-1$ as training data and made predictions for the image on trial $t$. For the category learning task, we used an $\ell_{2}$ regularised logistic regression model, and for the reward learning task, we used a Bayesian linear regression model with spherical Gaussian priors. We used the estimates from the linear models to predict participant choice using mixed-effects logistic regression in leave-one-trial-out cross-validation. For the category learning task, we regressed the probability estimates of the logistic regression models onto participant choice. For the reward learning task, we regressed the reward estimate differences between the left and the right options onto choice. Example learning curves for the two tasks are shown in Fig. 2C and Fig. 2D. Finally, we measure alignment to human choices using McFadden’s $R^{2}$ [57], which is computed as follows:
@@
Most of the representations we tested can do our task and predict human behaviour above chance level across the two tasks (as visualized in Fig. 3A and Fig. 3B). CLIP models were the top $7$ ($6$) models for the category (reward) learning task in predicting human choices. In total, $16$ ($7$) of the $86$ candidate representations predicted participant behaviour better than the ground truth representations that were used to generate the task. Of these $16$ ($7$) representations, $14$ ($6$) were CLIP models. One was a large vision transformer, trained in a supervised manner on ImageNet [58]. A human-aligned variant of DINO-v2 provided a better fit than the generative task representations in the category learning task. The rest of the supervised and self-supervised vision models, as well as the language models, had a heterogeneous distribution in how well they predicted human behaviour. To provide better intuition for how human participants and CLIP were similar, we display example trials where both CLIP and humans make the same incorrect decisions in Fig. 11 in Appendix C.
@@
Figure 5: Lower intrinsic dimensionality is linked with higher alignment most strongly for the multimodal models, and to a lesser extent with supervised ones.
@@
First, we found that none of the alignment methods improved alignment in our task consistently, with some instances of Harmonized and gLocal models improving alignment (Fig. 7A). Alignment improved task accuracy on average for Harmonized and gLocal models (Fig. 7B). Class separation was lower for all Harmonization and DreamSim models, whereas it increased for all gLocal models tested (Fig. 7C). We also observed that the similarity between the representations and the task embedding decreased after alignment for Harmonization and DreamSim, but it increased in most of the models after gLocal alignment (Fig. 7D). Lastly, we observed heterogeneous patterns in the change of intrinsic dimensionality across the three alignment methods, with gLocal reducing the intrinsic dimensionality for all but one of the tested models (Fig. 7E).
Figure 7: We compared models aligned to humans through three different methods against baselines that had the same architecture and that were pretrained on the same data. We found no consistent improvement in human alignment. Here plotted are (A) McFadden’s $R^{2}$, (B) task accuracy, (C) class-separation, (D) similarity with the task embedding, and (D) intrinsic dimensionality across model sizes and loss functions.
@@
Lastly, to better characterise how our cognitive tasks fit in the alignment literature, we compared them to previously established measures (Fig. 8). We found the strongest correlation with the THINGS odd-one-out judgements [3, 27] ($\rho=0.54$ for zero-shot, and $\rho=0.61$ for probing). Given the two tasks use the same images, and the ground truth of our tasks was constructed from the odd-one-out judgements, this strong relationship is expected. However, the correlations are still moderate, indicating important differences across tasks. Comparisons with an independent similarity judgement [23] task showed a weaker correlation ($\rho=0.35$), and we found no correlation with a fine-grained two-alternative forced choice task [12]. Lastly, we compared alignment in our task to alignment on the ClickMe dataset [76], which was used to build the Harmonization models [11]. We observed a negative correlation ($\rho=-0.48$) here, suggesting that pixel-level alignment and semantically bound global image alignment might be at odds.
@@ 5 Discussion
Figure 8: How do our tasks compare to other alignment methods? Our tasks offer similar (but not identical) results with two of the four similarity judgement tasks. There is a strong negative relationship with the ClickMe dataset, which focuses on localised pixel-level alignment.
@@ 5.1 Related work
Lastly, we found that a method designed for increasing human alignment, DreamSim[12], actually hurt alignment in our tasks. On the other hand, gLocal [10] and Harmonization [fel_harmonizing_2022] improved both performance and human alignment for some models but not all of them. However, the gLocal transform heavily utilises the THINGS dataset, as it made use of the triplet odd-one-out similarity judgement data [27], making it difficult to interpret how well it generalises to other settings. Taken together, these results highlight the importance of studying how well different alignment methods transfer across tasks, as we have done in this work.
@@ Appendix A Methods
Appendix B Testing aligned models on other datasets
Above, we also report how different alignment methods perform on different datasets(Fig. 9). Harmonization is on average slightly more human-like on the two external tested datasets compared to baselines. DreamSim shows mixed results for the Peterson et al. [23] dataset, but it shows improvement on the NIGHTS dataset [12]. This is not surprising, as this dataset was used to build DreamSim. Lastly, gLocal shows mixed results.
@@ Appendix B Testing aligned models on other datasets
Appendix C Additional results
Below we provide some additional results supporting our claims in the main text. Fig. 10 shows participants can do both tasks above chance level very early on in the task. Fig. 11 shows some incorrect choices made by humans and also by CLIP models, and Fig. 12 shows pairwise correlations between the factors we investigated that contribute to alignment.